Circular No. 6925

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998ck IN ESO 434-G20 J. Maza, Department of Astronomy, University of Chile, reports the

discovery by Marina Wischnjewsky, on a T-Max 400 film taken by L. E. Gonzalez with the Maksutov telescope at Cerro El Roble on May 31.969 UT, of an apparent supernova (B ∼ 17) located at α = 9^{h}41^{m}06^{s}.00, δ = −29^{o}06′17′′.7 (equinox 2000.0), which is 7′′ west and 1′′ south of the center of the spiral galaxy ESO 434-G20. The star was confirmed on direct CCD images obtained by I. Perez on June 1 using the 1.0-m telescope at Las Campanas Observatory, but it was not present on a film obtained on May 19.1 (B > 20).

SUPERNOVA 1998cg IN ANONYMOUS GALAXY F. Patat, European Southern Observatory (ESO); and M. Turatto,

Osservatorio Astronomico di Padova, observed SN 1998cg with the ESO 3.6-m telescope (+ EFOSC2; range 330–750 nm, resolution 1.8 nm) on May 30.07 UT at La Silla. The reduced spectrum shows that this object is indeed a supernova. The spectral features are consistent with a type-Ia classification, ∼ 1 month after maximum light (cf. Patat et al. 1996, MNRAS 278, 111). The spectrum is dominated by several emission lines with the usual P-Cyg profiles, superimposed on a red continuum. Most of the lines arise from Si II, S II, Fe II, and Fe III. The redshift of the parent galaxy, estimated from the Hα line, is 0.119, which confirms that the parent galaxy is a foreground object with respect to the cluster Abell 1514.

W COMAE BERENICES G. Tagliaferri and G. Ghisellini, Osservatorio Astronomico di Brera,

Milan; and P. Giommi, BeppoSAX, Rome, on behalf of the BeppoSAX ToO collaboration on blazars, report: “This BL Lac object (= ON 231) was observed with the BeppoSAX Narrow Field Instruments on May 11.96– 12.90 UT. The source was bright, with variability by a factor of ∼ 3 detected only below 2 keV. The source was well visible in the Phoswich Detector System up to ∼ 100 keV. The preliminary analysis indicates an average energy flux of 2.6 × 10^{−11} erg cm^{−2} s^{−1} in the band 0.1–4 keV. Above 4 keV, the spectrum hardens dramatically, with a 4–100-keV flux of ∼ 4 × 10^{−11} erg cm^{−2} s^{−1}. The extremely flat high-energy spectrum suggests a very high γ-ray flux. Continuing observations at all wavelengths are urged.”

1998 June 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green